OpenStudy (anonymous):

Bruce pitches for his little brother's baseball team. He has observed that the number of pitches a batter hits varies and is given by the function f(x) = x – 11, where x is the number of pitches and f(x) is the number of hits. If batters get {4, 12, 14, 27, 42} hits, how many pitches did Bruce throw? Select one of the options below as your answer: A. {13, 21, 24, 26, 44} B. {17, 24, 27, 28, 35} C. {20, 27, 32, 33, 45} D. {15, 23, 25, 38, 53}

OpenStudy (anonymous):

Substitute: {4, 12, 14, 27, 42} into f(x) = x – 11 as the f(x). I.e. 4 = x - 11, x=? solve for the rest

OpenStudy (anonymous):

d {15,23,25,38,53}